A rectangular garden is to be surrounded by a walkway of constant width. The garden's dimensions are 30 ft. by 40ft. The total area, garden plus walkway, is to be 1800ft^2.What must be the width of the walkway to the nearest thousandth?
A. 3.860ft
B. 2.610FT
C.5.430FT
D.7.720FT

Answer by venugopalramana(3286)

A rectangular garden is to be surrounded by a walkway of constant width.(X SAY) The garden's dimensions are 30 ft. by 40ft.
The total area, garden plus walkway, is to be 1800ft^2.
L OF GARDEN AND PATH =X+40+X=2X+40
W OF GARDEN AND PATH = X+30+X=2X+30
AREA OF GARDEN AND PATH = (2X+40)(2X+30)=1800
4X^2+140X+1200=1800
4X^2+140X-600=0
X^2+35X-150=0
X = [-35+ OR - SQRT(35^2+4*1*150)]/2 =3.86
A IS THE ANSWER
What must be the width of the walkway to the nearest thousandth?
A. 3.860ft
B. 2.610FT
C.5.430FT
D.7.720FT